Near the end of a marathon race, the first two runners are separated by a distance of 45.0 m. The front runner has a velocity of 3.45 m/s, and the second a velocity of 4.20 m/s.

(a) What is the velocity of the second runner relative to the first?
m/s faster than the front runner.
(b) If the front runner is 250 m from the finish line, who will win the race, assuming they run at constant velocity?

The first runner will win.The second runner will win.      


(c) What distance ahead will the winner be when she crosses the finish line?
m
Answer #1

here,

velocity of front runner , v1 = 3.45 m/s

velocity of second runner , v2 = 4.2 m/s

initial separation , s0 = 45 m

a)

the velocity of the second runner relative to the first , v = v2 - v1

v = 4.2 m/s - 3.45 m/s = 0.75 m/s

b)

distance of front runner from the front line , s = 250 m

let the time taken by runner to cross the finishing line be t1 and t2

for first runner

t1 = s /v1 = 250 /3.45 = 72.46 s

for the second runner

t2 = ( s+ s0) /v2 = ( 250 + 45) /4.2 = 70.24 s

as t2 < t1

the second runner will win

c)

the distance traveled by 1 in t2 time , s12 = v1 * t2

s12 = 3.45 * 70.24 m = 242.32 m

so, the distance ahead will the winner when crosses the line = s - s12 = 7.68 m